### v3

# Crop Pest Classification Using Deep Learning Techniques: A Review

Abstract
Insect pests continue to bring a serious threat to crop yields around the world, and traditional methods for monitoring them are often slow, manual, and difficult to scale. In recent years, deep learning has emerged as a powerful solution, with techniques like convolutional neural networks (CNNs), vision transformers (ViTs), and hybrid models gaining popularity for automating pest detection. This review looks at 37 carefully selected studies published between 2018 and 2025, all focused on AI-based pest classification. The selected research is organized by crop type, pest species, model architecture, dataset usage, and key technical challenges. The early studies relied heavily on CNNs but latest work is shifting toward hybrid and transformer-based models that deliver higher accuracy and better contextual understanding. Still, challenges like imbalanced datasets, difficulty in detecting small pests, limited generalizability, and deployment on edge devices remain significant hurdles. Overall, this review offers a structured overview of the field, highlights useful datasets, and outlines the key challenges and future directions for AI-based pest monitoring systems.
1 Introduction
Insects pests are one of the most persistent challenges in modern agriculture, with the global estimates indicating up to 40% crop yield loss annually due to pest [1]. Traditional pest monitoring, relying on manual inspections by farmers and entomologists, is labor-intensive, time consuming, subjective, and impractical for large-scale agriculture [2]. Recent advancements in deep learning, notably convolutional neural networks (CNNs) and vision transformers (ViT’s) , have shown great promise, achieving over 90% accuracy under controlled conditions for pest identification [3]. However, applying these technologies in real-world agricultural settings introduces complexities like variable lighting, occlusions, and complex backgrounds, significantly impacting performance [4].
The pest identification in agricultural fields has evolved rapidly from basic CNN architectures to sophisticated hybrid models that combine CNNs with vision transformers, leveraging the strengths of both local feature extraction and global context interpretation [5]. Recent innovations incorporating attention mechanisms and multi-scale processing have improved detection of small, camouflaged pests in complex environments [6]. Additionally, optimized mobile implementations now enable real-time pest identification via smartphone cameras, enhancing accessibility for farmers directly in the field [7]. Despite these technological advancements, dataset limitations persist, including severe class imbalance and insufficient representation of rare or region-specific pests, limiting model generalizability [8] [9]. Emerging methods such as quantum-inspired CNNs [10] and diffusion-based data augmentation [11] show potential but require further validation under practical agricultural conditions.
This review aims to provide a comprehensive analysis of research progress in deep learning-based crop pest classification, focusing on different types of crops and pests. The paper has four key contributions. First, we present a taxonomy of pest classification approaches in Section 3, organizing studies by pest types, target crops, techniques, datasets, and challenges. The studies based on different crops and pests are discussed in sections 4 and 5 respectively. Second, we provide detailed insights on technical approaches including CNN-based methods (Section 6.1), vision transformers (Section 6.2), hybrid architectures (Section 6.3), and object detection models (Section 6.4). Third, we analyze benchmark datasets and their characteristics in Section 7, identifying gaps in current data resources. Finally, it highlights important research gaps and open challenges that future studies need to address in Section 8, and presents the conclusion in Section 9.
2 Materials and Methods
To conduct this review, a structured and systematic search strategy was utilized to identify relevant studies on insect pest classification employing deep learning and artificial intelligence (AI) techniques. The search covered publications from approximately 2018 to 2025, reflecting the rapid advancements in AI-driven agricultural applications. Multiple scholarly databases, including IEEE Xplore, ACM Digital Library, SpringerLink, ScienceDirect, and Google Scholar, were searched using keywords such as ”insect pest detection,” ”pest classification,” ”deep learning of pests of crops,” ”plant disease and pest identification,” ”YOLO pest detection,” and ”vision transformer agriculture.” Initial searches identified 355 articles. After removing duplicate entries (n = 21), 334 records remained. These records underwent a screening process based on their titles and abstracts, which led to the exclusion of 172 irrelevant articles, leaving 166 for further review. Subsequently, full-text retrieval was conducted for these 166 articles. All reports were successfully retrieved and reviewed. The inclusion criteria applied during full-text assessment required that (1) studies focused explicitly on insect pests affecting crops (excluding studies exclusively addressing plant diseases or weeds), (2) methods involved AI or machine learning with computer vision, (3) publications were peer-reviewed journal or conference papers or reputable preprints, and (4) studies provided adequate technical details and experimental results for evaluation.
After the full-text assessment, 121 reports were excluded: 51 studies were not focused specifically on insect pests, 43 lacked detailed technical methodology or sufficient experimental results, and 27 did not employ AI or ML-based image analysis for classification. This rigorous selection process resulted in a final set of 37 studies included in this review. Throughout the process, efforts were made to ensure comprehensive coverage, including foundational early work in convolutional neural networks (CNNs) and recent advancements such as transformer models. Additionally, reference lists of key selected articles were reviewed to include significant studies potentially overlooked during the database search. The final collection of papers encompasses a global body of work, including several surveys that contextualize the field (e.g., Teixeira et al., [9]) as well as diverse case studies on pest detection in different crops and environments. This systematic approach to literature gathering ensures that the insights drawn in this chapter are representative of the current state of research in AI-based pest classification. The figure 1 illustrates the process used for identifying, screening, and including studies in this review.
Fig 1: PRISMA flow diagram for literature selection.
3 Taxonomy of AI Based Pest Classification
This section presents a comprehensive taxonomy developed from the reviewed literature on AI-assisted pest classification. The taxonomy is organized around five key dimensions: the types of insect pests studied, the crops targeted, the AI techniques employed, the challenges encountered, and the datasets used. This structured classification allows us to analyze trends across research works and highlight underexplored combinations, such as underrepresented pest species or limited use of hybrid attention-based models like HPMA-ViT. The following subsections present each of these dimensions with supporting literature, tables, and insights. This categorization provides a structured view of the research landscape and will be accompanied by a detailed explanation of each category in the subsections below. The figure 2 illustrates the categorization of pest types, associated crops, AI techniques, benchmark datasets, and key challenges addressed in recent literature.
Fig 2: Taxonomy of pest classification approaches.
4 Crops Addressed
Research on AI-driven pest classification has spanned a variety of crops, often focusing on those of high economic importance or with well-documented pest issues. We organize the literature by the main crops addressed, as different crops present different pest challenges and imaging conditions. Table LABEL:tab:ai-crop summarizes representative studies for each major crop, including the reference, year, brief description, and dataset used.
4.1 Rice
Rice is a staple crop worldwide and hosts numerous insect pests (planthoppers, stem borers, leaf folders, etc.). Several studies have targeted rice pest identification. For example, Liu et al. [12] developed a CNN-based system to classify 12 common paddy field insect pests, using a dataset of 5,000 images covering leafhoppers, brown planthoppers, and other rice pests. Their approach achieved promising accuracy but was limited by relatively few samples per class (428 images on average), reflecting the data scarcity at the time. More recently, object detection models have been applied to rice pests. Du et al. [13] introduced an improved YOLOv7 model (named YOLOv7-PSAFP) to detect small pests in rice and corn fields. By enhancing the feature pyramid and loss functions to better capture tiny objects, their model exceeded 93% mean Average Precision (mAP) on a custom corn-and-rice pest test set. This was a significant improvement over baseline detectors (the standard YOLOv7 and Faster R-CNN) on the same dataset. Another example is the work by Javeria Amin et al. [10], who combined YOLOv5 for rice pest localization with a quantum convolutional network for classification. They treated it as a binary problem (paddy with pest vs. without pest) and reported extremely high accuracy (99.9%) on their paddy field images. These rice-focused studies highlight how both CNN classifiers and one-stage detectors have been customized for the rice domain. However, multi-class rice pest identification (distinguishing among many pest species in rice) remains challenging unless large labeled datasets are available. The IP102 benchmark, which includes many rice pest species, has exposed the difficulty of fine-grained classification with a baseline accuracy below 50% (ResNet-50). Nonetheless, progress is steadily being made via improved models and data augmentation targeted at rice pest images.
4.2 Tomato
Tomato crops also feature prominently in pest classification research due to their economic value and susceptibility to pests (whiteflies, caterpillars, etc.). Al-Shannaq et al. [3] (2025) focused on tomato pest infestation identification using a basic CNN model augmented with extensive image preprocessing. They assembled a dataset of tomato leaves showing pest damage and boosted its size with image augmentation techniques. The CNN achieved about 85–90% accuracy on distinguishing pest-damaged vs. healthy tomato leaves. Notably, without augmentation, the performance dropped to nearly 50–60%, underscoring how limited data can hurt generalization. This study’s outcomes indicate that augmentation was critical to reach high accuracy on the small tomato dataset, and it suggested the need for larger, more diverse training data. In another tomato-related work, Liu and Wang [14] applied the YOLOv3 object detection model (with multi-scale image pyramids) to detect tomato pests and diseases in images. Their improved YOLOv3 could identify multiple issues on tomato plants (like pest presence and disease symptoms) with high precision, although exact performance metrics were not given in the excerpt here. Challenges noted in tomato pest studies include dealing with complex backgrounds (foliage, varying lighting) that can confuse classifiers. For instance, Ramcharan et al. [4] achieved 93% accuracy on field images of cassava diseases but observed some misclassifications due to lighting and background noise, a problem likewise relevant to field images of tomato pests. Overall, the literature on tomatoes shows that CNNs and YOLO detectors can be effective for targeted pest problems, provided the models are trained on sufficiently representative images or enhanced via transfer learning and augmentation. Table LABEL:tab:ai-crop lists these tomato-focused studies, noting their datasets (often custom images from greenhouses or fields) and results.
4.3 Groundnut
Peanut crops (groundnuts) have been the subject of pest classification research, especially in recent years with the advent of transformer models. Two independent studies by Venkatasaichandrakanth and Iyapparaja [15] targeted a peanut pest identification using a hybrid CNN-transformer architectures. Nagalakshmi et al. [16] proposed an Enhanced Vision Transformer Architecture (EViTA) tailored for peanut pests. Their model uses dual encoders: a CNN backbone for local feature extraction and a transformer module for global context, combined with a Moth Flame Optimization (MFO) algorithm for feature selection. Evaluated on a peanut pest image dataset, EViTA achieved about 92% classification accuracy and a precision of 0.95 in distinguishing multiple peanut pest species. Similarly, Venkatasaichandrakanth and Iyapparaja [15] developed a dual-layer transformer model (also incorporating MFO) for peanut pests, reporting virtually the same accuracy (92%) and even a perfect specificity on their test set. The convergence of these two studies in approach and performance suggests a strong interest in transformer-based solutions for pests in groundnut fields. These models demonstrated that incorporating attention mechanisms can improve accuracy over traditional CNN-only models for the fine-grained task of identifying peanut pests. The peanut pest studies also noted that the hybrid models were particularly good at integrating multi-scale features (leaf damage patterns, pest body details), which is important in cases where pests vary in size from tiny aphids to larger caterpillars attacking peanut plants. The datasets used in these works were relatively small (on the order of a few thousand images), often collected from local peanut fields, so results might be optimistic for those specific conditions. Nonetheless, the peanut case exemplifies how researchers are pushing beyond CNNs to boost pest classification performance in specific crops.
4.4 Tea
Tea plantations face unique pest challenges, including very small insects like mites and leafhoppers that can be hard to spot. Recent AI research on tea crops has emphasized detecting tiny pests on tea leaves. For instance, Zejun Wang et al. [17] proposed a lightweight YOLOv8-based model to detect microscopic pests in tea gardens. By introducing attention mechanisms and adaptive convolution modules, their improved YOLOv8 achieved significantly higher mAP (mean Average Precision) than the baseline increasing by 5–16% in some metrics, enabling fast and accurate identification of early-stage tea pests. This indicates the model could reliably spot small sap-sucking pests (like tea green leafhoppers or mites) that often appear as only a few pixels in images. Earlier work by Samanta et al. [18] had used traditional feature selection and neural networks on a much smaller tea pest dataset (8 species, 609 images), but modern deep learning greatly surpasses those results. Another study by Wang et al.[19] used a sliding window CNN approach to detect tea pests in images, illustrating an alternate method to handle tiny insects by brute-force scanning. Across these studies, tea emerges as a crop where high image resolution and specialized models are needed to overcome background clutter (leaf textures) and the minute size of pests. The improved results with attention-enhanced YOLO models (e.g., adding spatial attention to focus on pest regions) show that deep learning can indeed tackle even the small-pest problem in tea. Table LABEL:tab:ai-crop compiles the efforts in the tea domain, including the datasets (often custom-collected images of tea leaves in field conditions) and techniques ranging from classical to cutting-edge.
4.5 Cactus
An unusual but noteworthy case is pest management in cactus cultivation. Berka et al. [7] introduced CactiViT, a vision transformer-based system to detect infestations of the carmine cochineal (a scale insect) on prickly pear cacti. This pest is devastating in regions like Morocco and requires early intervention. The authors created a new image dataset of cochineal-infested cacti and developed a mobile application powered by a ViT model to classify infection levels. Their transformer model (ViT-B16 backbone) achieved 88.7% classification accuracy on the cactus infestation dataset, outperforming comparable CNN models by about 2.6%. Impressively, the system runs in real-time on smartphones, allowing farmers to photograph a cactus and immediately receive an assessment of pest presence and severity. This study demonstrates the viability of transformers even on relatively small, custom datasets – likely aided by transfer learning and the strong representational capacity of ViTs. It also highlights how AI can be applied beyond staple crops, tailored to specific regional pests. The CactiViT dataset has been made publicly available, which is valuable given the lack of open data on this pest. While cactus farming is a niche compared to rice or maize, the success of CactiViT underlines that AI-assisted pest management principles extend to any crop: with a suitable dataset and model, even a challenging pest like cochineal (which blends into its host plant) can be reliably detected. This example is included in our review as it pushes the envelope on deploying AI (transformers on mobile devices) for pest monitoring in the field.
4.6 Other Crops
In addition to the above, many studies cover pests in other crops such as maize, cassava, cotton, and orchard trees. For instance, cassava pest and disease recognition was explored by Lin et al.[20] using a Graph Pyramid Attention Network. While they achieved near 99% accuracy on a confined cassava disease dataset, the performance dropped to 57% when the same model was tested on the diverse IP102 insect dataset. This suggests that models trained on one crop/pest scenario (cassava) may not generalize well to others without retraining. Maize (corn) pests like fall armyworm have been targeted by object detectors; researchers have developed optimized YOLO models for corn fields, as mentioned by Du et al.[13] for corn and rice pests. Cotton pests and soybean pests have likewise seen deep learning solutions, though they are fewer in number. Some recent works tackled orchard pests, for example, Liu et al.[12] designed a transformer-based detector for forestry pests (e.g., insects on tree leaves) and reported 95.3% mAP on a constrained forestry dataset, though this plummeted to 35.6% mAP on the open-domain IP102 set. Stored grain pests (like beetles in stored products) have also been studied with CNN classifiers, often using Xie’s datasets of common storage pests. In summary, beyond the major crops, AI techniques have been applied to a wide range of agricultural contexts. However, the success in one domain does not always translate to another, due to differences in pest appearance, background, and image capture conditions. Table LABEL:tab:ai-crop encompasses a selection of studies across various “other” crops, illustrating the breadth of applications and reminding us that each crop-pest combination may need specific consideration in model design and training.
Table 1: Summary of deep learning techniques for crop pest/disease detection.
Ref (Year) Crop AI Technique Dataset (Images, Classes) Results Notes/Limits
Mohanty et al.[2] Multiple (14 crops) CNN (AlexNet, GoogLeNet) PlantVillage (54k images, 26 diseases) 99.3% acc (lab) High accuracy on lab images; performance drops in field.
Ramcharan et al.[4] Cassava CNN (ResNet-50, transfer) 5,660 field images (5 classes) $\sim$93% acc (field) Field conditions cause some errors (lighting, background).
Wu et al.[21] Various (8 crops) CNN (ResNet-50 baseline) IP102 (75k images, 102 pests) 49.4% acc (baseline) Large, imbalanced dataset; fine-grained classes difficult.
Al-Shannaq et al.[3] Tomato CNN (custom + augmentation) 5,894 tomato leaf images (augmented) 85% test acc Small original dataset ($\sim$859 images); augmentation essential.
Zhang et al.[17] Tomato Object Detection (YOLOv3+) Tomato pest/disease images (custom) High mAP (not stated) Improved multi-scale detection on tomato plants; field validation needed.
Amin et al.[10] Rice/Paddy YOLOv5 + Quantum CNN (hybrid) Paddy Pest (Kaggle: $\sim$648 images, binary) 99.9% acc (binary) Innovative quantum approach; only binary classification (pest vs no pest).
Li et al.[22] Multi-crop Vision Transformer (ViT) 39-class PlantVillage (pest & disease) 96.7% acc Outperformed CNNs on same data; lab images, needs field testing.
Berka et al.[7] Cactus Vision Transformer (ViT-B16) CactiViT dataset ($\sim$1k images, 3 classes) 88.7% acc Mobile real-time app; dataset small but open.
Nagalakshmi et al. [16] Peanut Hybrid CNN+ViT (EViTA + MFO) Peanut pest images ($\sim$2k, 4 classes) $\sim$92% acc, 0.95 precision Improved over CNN; custom dataset, specific to region.
Venkatasaichandrakanth [15] Peanut Hybrid (dual-transformer + MFO) Peanut pest images ($\sim$2k, similar to above) $\sim$92% acc, high specificity Similar approach to [16] confirms transformer benefit.
Xiao et al.[23] Multi-crop Efficient CNN (GA-GhostNet) IP102 (102 classes); others via transfer 71.9% acc (IP102); 95% on simpler Lightweight (4M params) model; struggled on large 102-class task.
Du et al.[13] Corn/Rice Object Detection (YOLOv7-PSAFP) Corn-rice pest dataset (custom, 20+ pests) 93% mAP (field) Optimized for small pests in clutter; outperforms std. YOLOv7.
Wang et al.[6] General Object Detection (YOLOv8+CBAM) Insect-YOLO (3M param model, custom data) 93.8% mAP (field) Tiny model for edge devices; effective on low-res drone images.
Kar et al.[24] Multi-crop Self-Supervised (SSL ResNet) Unlabeled + labeled pest images 93–94% acc (with fewer labels) SSL pre-training reduced label needs; not widely adopted yet.
5 Pests
Sap-sucking Insects
The insect pests addressed across the studies can be grouped into broad types. One major category is sap-sucking insects (order Hemiptera), which includes aphids, planthoppers, leafhoppers, whiteflies, and related pests. These are small insects that feed on plant sap and often occur in large numbers. They appear in multiple studies: for instance, aphids are one of the target classes in peanut and groundnut pest works ( Venkatasaichandrakanth et al.[15], as well as in the GA-GhostNet evaluation (Xiao et al.[23], used the IP102 dataset where aphids are a category). Planthoppers (e.g., the brown planthopper) are serious rice pests – Wang et al. explicitly mention rice planthopper and report improved precision for it using their Insect-YOLO model. Sap-sucking pests present challenges because of their small size; as noted by Wang et al., baseline detectors had low precision on tiny planthoppers until attention mechanisms were added. These insects also often cause indirect damage by transmitting viruses, so identifying their presence early is critical.
5.1 Caterpillars and Borers
Another important group is caterpillars and borers, which are larvae of moths and butterflies (order Lepidoptera). These are chewing pests that consume leaves or bore into stems. Many studies include examples: Hu et al.[25] dealt with rice pests like the rice leaf roller and stem borers (both are caterpillar stages of moths). The peanut “Gram Caterpillar” (likely referring to Helicoverpa or a similar genus) is included in Nagalakshmi et al. [16]. Armyworms – a type of noctuid moth caterpillar – appear in the groundnut pest classification (Venkatasaichandrakanth et al.[15]) and in the IP102 dataset used by others. Caterpillars tend to create visible damage (holes in leaves, defoliation), which can aid detection; however, they can also hide (e.g., inside stems for borers), which is an ongoing challenge. Some models, like Liu et al.[12], aimed to improve detection of occluded or small pests which could include larvae that are hidden or early instars that are tiny.
5.2 Beetles and Weevils
Beetles and weevils (order Coleoptera) form another pest type covered in some datasets. For instance, the IP102 dataset (used in Lin et al [20] and others) contains various beetle species (e.g., flea beetles, weevils). The GA-GhostNet paper (Xiao et al.[23]) mentions that their model was tested on the jute pest dataset, which includes jute hairy caterpillar and yellow mites, and also on an apple dataset (which might involve codling moth larvae or weevils). While specific beetle pests were not individually highlighted in many deep learning studies (perhaps due to fewer public image sets), the category remains important – e.g., the Colorado potato beetle, various weevils in stored grains, etc., are potential targets for future AI models. One reason beetles may be less rep- resented so far could be that some beetles are best detected with traps rather than imagery (e.g., pheromone traps for weevils).
5.3 Other Insects
Lastly,this category includes miscellaneous or less common pest types like flies (e.g., fruit flies), wasps, thrips, or any insect pests not in the above categories. Thrips (tiny fringe-winged insects) were a class in the groundnut pest study (thrips of-ten attack peanuts and soybeans by sucking plant juices similar to aphids). Fruit flies or maggots were not directly addressed in the reviewed papers, but they are candidates for similar image classification techniques (perhaps via trap images). In general, the reviewed works collectively cover a wide taxonomic range – from soft-bodied aphids to hard-shelled beetles and moth larvae – demonstrating the versatility of AI mod- els. Each type poses its own detection challenges: e.g., aphids require high-resolution imagery to resolve their tiny bodies, moths/caterpillars require temporal monitoring since they grow quickly, and borers might require detecting indirect symptoms (like frass or entry holes) rather than the insect itself. The table 2 highlights these groupings, and indeed many studies implicitly focus on one group or another (for example, the tea and cactus studies effectively dealt with pest damage symptoms, which is common for sap-suckers and disease-causing insects).
Table 2: Recent image-based AI studies for insect pest types, datasets, and challenges.
Ref Pest Type Examples (Species) Datasets / Crops Remarks
[17], [7], [13] Sap-Sucking Insects Aphids, Whiteflies, Leafhoppers, Scales IP102 (many aphid/whitefly classes); Tea pest images (custom); CactiViT dataset (cactus) Tiny size; require high-res and attention. Class imbalance common (few samples of certain species).
[26],[2],[13] Caterpillars & Borers Fall armyworm, Stem borers, Leaf miner larvae PlantVillage (includes some pest damage); Custom field images (maize, tomato) Larger larvae easier to classify if visible. Hidden borers still a challenge (may rely on damage signs).
[27],[28],[29] Beetles & Weevils Grain weevils, Colorado potato beetle, Palm weevil Xie et al. 2018 (40 pest dataset); NBAIR (India) pest images; Custom trap images Many lab datasets exist (isolated insect images). Field detection requires dealing with camouflage and varied backgrounds.
[30], [31] Other (Flies, etc.) Fruit flies, Locusts, Thrips, Mites, etc. Custom trap image sets; Drone imagery (locusts); Mixed insect image datasets Highly variable group. Large-scale models struggle across all (e.g., IP102). Few-shot learning emerging to handle rare pests.
6 Techniques and Approaches
6.1 CNN Based Approaches
Convolutional Neural Networks (CNNs) have been the workhorse of image classification and thus dominate early pest identification literature. These models excel at capturing local spatial features via convolutional filters. In pest detection, many studies leveraged classic CNN architectures pre-trained on ImageNet (e.g., VGG, ResNet) and then fine-tuned on pest datasets. For example, Ayan et al.[32] tested multiple CNN architectures (VGG16, ResNet50, InceptionV3, Xception, MobileNet, etc.) on crop pest images, achieving around 97% accuracy with top models. Similarly, Al-Shannaq et al.[3] evaluated GoogleNet (a CNN) as a baseline in their pest classification framework, illustrating the strong performance of CNNs on insect image data. To improve upon standard CNNs, researchers proposed specialized or optimized CNN variants. Sandhya Devi et al.[33] employed an EfficientNetV2 model (a CNN with compound scaling) to handle the high variability in field-captured pest images, reporting improved accuracy and robustness for classifying crop pests. In another notable work, Xiao et al.[23] introduced a custom lightweight CNN called GA-GhostNet, which integrates attention modules to drastically reduce model size without sacrificing accuracy. Their GhostNetbased classifier retained performance while being compact enough for deployment on resource-constrained. In a related study, Marinho et al.[34] applied CNNs to X-ray imagery of fruit fly pupae, achieving over 97% accuracy in distinguishing parasitized versus healthy pupae – underscoring CNNs’ capability in even unconventional imaging modalities. Despite their effectiveness, CNNs have limitations when it comes to modeling global context or extremely fine-grained differences across an image. Pest species often require distinguishing subtle textures or shapes, and CNNs focusing on local receptive fields can miss these global patterns. These shortcomings of CNNs have paved the way for transformer-based architectures, which we discuss next (Dosovitskiy et al.[5]). Nonetheless, CNN-based solutions remain highly competitive for many classification tasks – often exceeding 90% accuracy on curated pest image dataset especially when sufficient training data is available and the model is carefully tuned. Table 3 present the summary of the Key recent CNN based approaches for crop pest and disease detection.
Table 3: Literature Review Table on CNN-Based Approaches.
Ref Year Description
[32] 2020 Evaluated 7 CNN architectures (VGG16, ResNet50, InceptionV3, etc.) for multi-pest classification; top models achieved $\sim$97% accuracy.
[3] 2025 Used GoogleNet (CNN) as a baseline for pest ID in Jordan; highlighted CNNs’ strong performance on local pest datasets.
[33] 2023 Employed EfficientNetV2 CNN to handle high variability in field images, improving pest classification accuracy.
[23] 2024 Proposed GA-GhostNet (a lightweight CNN with attention) achieving competitive accuracy with dramatically fewer parameters.
[34] 2023 Applied CNN models to X-ray images of fruit fly pupae; achieved >>97% accuracy in classifying parasitized vs. healthy pupae.
6.2 Vision Transformers
Transformers originally popularized in NLP have emerged as powerful vision models due to their ability to capture global context via self-attention. Vision Transformers (ViTs), first introduced by Dosovitskiy and colleagues [5], treat an image as a sequence of patches and learn long-range relationships, overcoming the locality bias of CNNs. Starting around 2020–2021, ViTs began to permeate pest classification. These models show particular value for high-resolution images or scenarios requiring holistic image understanding. In pest detection, Li et al.[22] demonstrated that a ViT model, when sufficiently trained and augmented, could achieve up to 96% accuracy on a mixed pest and plant disease dataset – outperforming comparable CNNs on the same data. This highlights transformers’ advantage in modeling global image features (such as an insect’s overall shape or the context of its surroundings) which are critical for distinguishing certain pest species. Similarly, Fu et al.[35] improved ViTs’ ability to focus on important regions by introducing a block-wise image partitioning before the self-attention layers. This modification effectively guided the transformer to attend to biologically significant areas (e.g., lesion spots or insect body parts) and yielded higher accuracy in pest identification. Transformers have also been integrated into object detection frameworks for pests. For example, Hu et al.[25] embedded a Swin Transformer module into a YOLO detector’s feature pyramid. The resulting model (sometimes termed “YOLO-GBS”) captured long-range dependencies in complex rice field images, improving pest detection performance under cluttered backgrounds. In another work, Liu et al.[12] developed a fully transformer-based detector for pests, incorporating a specialized Region-based Pyramid Self-Attention (RPSA) mechanism to better detect small insects and distant. These studies demonstrate how transformers can overcome some CNN limitations – especially in global reasoning and nuanced differentiation among visually similar classes. It is worth noting that ViTs typically require large training datasets to realize their full potential, due to fewer built-in inductive biases than CNNs. In the agricultural domain, where data can be scarce, techniques like data augmentation and transfer learning are often employed alongside ViTs (Li et al.[22]. Nonetheless, as open datasets grow and transformer architectures are refined, ViTs have begun to surpass CNNs in pest classification tasks that benefit from global feature modeling. The transformer approach marks a shift towards models that can consider an entire scene (e.g., a leaf and its pests) collectively rather than piecemeal.Table 4 illustrate the key studies leveraging vision transformers for pest and disease detection in agricultural imagery.
Table 4: Literature Review Table on Vision Transformer Based Approaches.
Reference Year Description
[5] 2021 Introduced Vision Transformer (ViT) for image recognition, using self-attention on image patches instead of convolutions, a foundational work enabling transformers in vision.
[22] 2022 Applied a ViT to pest images with extensive augmentation; achieved $\sim$96% accuracy, outperforming CNN baselines on the same pest/disease dataset.
[35] 2024 Improved ViT’s focus via block-wise image partitioning, helping the model attend to key pest regions (e.g., insect body), which boosted classification accuracy.
[25] 2023 Embedded a Swin Transformer into YOLO (YOLO-GBS) for rice pest detection; captured long-range dependencies in cluttered field images, improving detection performance.
[12] 2025 Developed a fully transformer-based pest detector with RPSA attention; excelled at modeling global context and detecting small pests in images.
6.3 Hybrid CNN-Transformer Architectures
To capitalize on both local and global feature learning, researchers have increasingly turned to hybrid models that combine CNN and transformer components. These architectures aim to harness CNNs’ prowess in capturing fine details (edges, textures) and transformers’ strength in modeling long-range relationships. A number of studies report that such hybrids outperform either pure CNNs or pure transformers in isolation. One exemplar is the EViTA model proposed by Nagalakshmi et al.[16], later extended by Iyapparaja [15]. EViTA (Enhanced Vision Transformer Architecture) uses CNN layers to first extract low-level features from pest images, which are then fed into a transformer encoder. This design provided a powerful blend of detailed texture representation and global context understanding. In tests on peanut crop pest images, the hybrid EViTA achieved higher accuracy than either a CNN-only or transformer-only approach, validating the complementary nature of CNN and transformer features (Nagalakshmi et al.[16]. Similarly, Zhan et al.[36] developed a hybrid model for tea pest detection: initial CNN layers captured local insect features, followed by iterative transformer blocks to integrate global information like surrounding leaf. Their hybrid significantly outperformed a baseline CNN when dealing with diverse backgrounds in tea fields. Another innovation by Saranya et al.[37] was to embed a hybrid pooled multihead attention (HP-MHA) module within a ViT framework. This HP-MHA module effectively fused CNN-style feature pooling with transformer attention. By doing so, the model retained fine-grained detail (through pooling operations) while still leveraging self-attention to capture global patterns. The result was improved accuracy on pest classification, especially for tasks requiring both detail sensitivity and contextual reasoning. In essence, Saranya et al.’s work moves toward hybrid attention, where elements of CNN architecture (like pooling) are integrated into transformer attention layers. Hybrids have also made their mark in object detection. Transformer-enhanced YOLO variants illustrate this trend: for instance, the aforementioned YOLO-GBS by Hu et al.[25] and the transformer-based detector by Liu et al.[12] can be viewed as hybrid architectures (CNN backbone + transformer neck). These systems showed improved detection accuracy and spatial precision, confirming that mixing convolutional and attention mechanisms can yield superior results for locating pests in images. Recently, Utku et al.[38] introduced ConvViT, a hybrid model for farm insect detection that combines a CNN feature extractor with a ViT classifier. ConvViT was optimized for edge devices, demonstrating strong performance on-device by balancing accuracy and efficiency. Likewise, Fang et al.[11] proposed Pest-ConFormer, a CNN-transformer hybrid with multi-scale feature selection, which achieved state-of-the-art results on a large-scale crop pest dataset (14 pest species) . Overall, hybrid architectures underscore the complementary nature of CNNs and transformers: CNN components excel at recognizing texture or fine details like wing venation or spot patterns, whereas transformer components capture the broader context such as pest position on a leaf or co-occurrence of multiple pests in one image. The synergy of these aspects leads to more robust models, particularly valuable in the complex scenes of agricultural settings. The table 5 lists the overview of recent hybrid-CNN approaches for crop pest and disease detection.
Table 5: Literature Review Table on Hybrid CNN–Transformer Based Approaches.
Reference Year Description
[16] 2024 Developed EViTA, a hybrid model with CNN feature extractor + transformer encoder, outperforming standalone CNN or ViT on peanut pest images.
[15] 2023 Extended EViTA architecture; integrated Moth Flame Optimization (MFO) for feature selection and parameter tuning, further enhancing pest classification precision.
[36] 2023 Hybrid CNN-transformer for tea pests: CNN layers for local features, followed by transformer blocks for global context—improved performance in cluttered field images.
[37] 2024 Embedded a hybrid pooled multi-head attention module into ViT, fusing CNN pooling with self-attention; achieved higher accuracy on fine-grained pest recognition.
[38] 2025 Introduced ConvViT for insect detection, a CNN+ViT hybrid optimized for mobile deployment, combining high accuracy with low model size (edge-friendly).
[11] 2024 Proposed Pest-ConFormer, a multi-scale hybrid CNN-transformer architecture for large-scale pest recognition, achieving state-of-the-art results on a 14-class pest dataset.
6.4 Object Detection Models
While classification (assigning a single label to an image) is useful, practical field scenarios often involve multiple pests per image or a need to pinpoint pest locations. Thus, object detection models have been applied to agricultural pest monitoring, with frameworks like Faster R-CNN and especially YOLO becoming dominant. In recent years, YOLO-based approaches are favored for their real-time performance and high accuracy trade-off. Amin et al.[10] utilized YOLOv5 to localize pests in field images, illustrating how detection networks can both identify and locate insects within an image frame. Building on this, Du et al.[13] and Wang et al.[17] experimented with newer YOLO versions (YOLOv7, YOLOv8) and reported improved detection accuracy and speed thanks to architectural advancements in those versions. For instance, Wang et al. integrated YOLOv8 into a pest monitoring system and even performed a regression analysis on the detection outputs to estimate pest population counts, achieving an R² of 0.99 compared to manual counting. Such a high correlation demonstrates that detection models are reliable enough for quantitative pest monitoring, a key requirement for Integrated Pest Management (IPM) programs. On the other hand, Faster R-CNN has served as a strong two-stage detector baseline. Teixeira et al.[9] noted in their systematic review that many early studies used Faster R-CNN for insect detection in traps, although YOLO’s one-stage approach often achieves comparable accuracy with significantly faster inference. In one example, Teixeira et al.[9] compared detectors and found YOLO models consistently outperformed others on the IP102 pest dataset (which contains 102 pest species) in both precision and speed. Consequently, recent work in pest detection overwhelmingly favors YOLO variants or similar one-stage architectures. Researchers have introduced domainspecific enhancements to these models. Zhou et al.[39] developed an Insect-YOLO model tailored to capturing small pests on various crops, featuring streamlined parameters for speed and custom data augmentation for diversity. This specialization yielded exceptional accuracy with real-time speeds, addressing some challenges of deploying detectors in the field. Similarly, Yu and Zhang [40] proposed a YOLOv5- based detector for rice pests with an improved feature pyramid to better detect tiny insects, significantly boosting recall for small pest targets. These targeted improvements show that with careful adaptation, general-purpose models like YOLO can be tuned to agricultural needs, e.g. focusing on small object detection or handling dense pest clusters on traps. An important aspect of detection in agriculture is dealing with class imbalance (many pest instances might belong to a few common species, with others being rare). Researchers often incorporate techniques like hard example mining or focal loss into detectors to bias training towards those hard-to-detect, less frequent pests. For instance, Du et al.[13] emphasized training on hard examples (e.g. pests camouflaged against leaves) to reduce background false positives. Such strategies improved robustness in complex field imagery. In summary, object detection models particularly the YOLO family are crucial for advancing from simply saying “this image has pest X” to answering “where and how many pests are present.” They enable automated pest counting and mapping, which are invaluable for precision agriculture. The literature trend clearly shows YOLObased methods leading the pack for pest detection, owing to their speed and adaptability, with continual refinements to address domain-specific challenges like tiny object size and cluttered backgrounds. Table 6 summarize the recent object detection models for agricultural pests detection and monitoring.
Table 6: Literature Review Table on Object Detection Models.
Reference Year Description
[10] 2023 Employed YOLOv5 for pest localization in field images; demonstrated one-stage detectors can accurately detect multiple pests per image.
[13] 2024 Applied YOLOv7 to pest detection with improved accuracy; utilized hard example mining to better learn from challenging, camouflaged pest instances.
[6] 2025 Deployed YOLOv8 in a pest monitoring system; achieved $R^{2}=0.99$ for automated pest counts, validating detection models for precise population estimation.
[9] 2023 Systematic review of insect detection; noted YOLO models consistently outperform Faster R-CNN on pest datasets in both accuracy and speed.
[40] 2023 Developed a lightweight YOLOv5-based detector for rice pests; added a large-scale feature layer to improve detection of small pests, boosting recall significantly.
[39] 2023 Created “Insect-YOLO” with custom enhancements for multi-crop pest detection (streamlined parameters, specialized augmentations); achieved real-time detection with high accuracy.
6.5 Emerging and Non-Traditional AI Techniques
Beyond mainstream deep learning approaches, a small but growing number of studies are exploring novel AI paradigms. Amin et al.[10] ventured into quantum machine learning, applying a quantum CNN to a pest classification task. Although the model was only tested on a small dataset, it achieved near-perfect accuracy, showcasing intriguing early potential for quantum-enhanced image processing albeit in a nascent, experimental stage More practical in the short term are self-supervised learning methods. Kar et al.[24]utilized BYOL (Bootstrap Your Own Latent) to pre-train pest classification mod- els using unlabeled data. Their approach nearly matched the performance of fully supervised counterparts, highlighting a valuable strategy for overcoming the labeling bottleneck in agricultural datasets. Additionally, interdisciplinary methods such as metaheuristic optimization have been applied. For instance, Moth Flame Optimization (MFO) was used by Nagalakshmi [16] and Venkatasaichandrakanth [15] to enhance feature selection and model parameter tuning. These methods, while less conventional in computer vision, demonstrate potential for improving performance, particularly in small-sample or constrained-data settings. Table 7 summarizes the novel and non traditional AI methods recently explored for crop pest detection and decision support.
Table 7: Literature Review Table on Emerging and Non-Traditional Approaches.
Reference Year Description
[10] 2023 Explored a quantum convolutional neural network for pest classification; achieved near-perfect accuracy on a small dataset, showing early promise for quantum-enhanced learning.
[24] 2022 Used self-supervised BYOL to pre-train on unlabeled pest images; fine-tuned model reached $\sim$94% accuracy on 12-class pest identification, nearly matching fully supervised results.
[16] 2024 Incorporated Moth Flame Optimization (MFO) in their hybrid model’s training; this metaheuristic optimized feature selection and hyperparameters, boosting classification performance.
[41] 2022 Applied a diffusion model to generate synthetic pest images for data augmentation; improved long-tail pest class accuracy significantly by augmenting scarce classes.
[42] 2022 Developed a GAN-based augmentation tool for pest detection (case study: whiteflies on traps); augmented data led to higher detection rates for rare pest instances.
[43] 2022 Integrated pest CNN models with an Augmented Reality app for farmers; enabled real-time pest identification through smartphone camera with overlaid guidance (enhancing practical usability).
The technical landscape of AI in pest classification is rapidly evolving from foundational CNN classifiers to transformer-centric and hybrid architectures, and even experimental paradigms like quantum learning and self-supervised approaches. The literature consistently shows that using advanced architectures such as ViTs, attention modules, or hybrid combinations yields significant performance improvements. However, such gains are often balanced against increased model complexity and resource demands. Authors like Saranya et al.[37] have explicitly noted the trade-offs between accuracy and deployability. As a result, there is growing interest in designing lightweight yet effective models. For example, Xiao et al.[23] and Wang et al.[6] focused on efficiency through model compression or architectural streamlining to enable on-device deployment, an important consideration for field applications where compute resources may be limited. Ultimately, the diversity of approaches provides researchers and practitioners with a robust toolkit. The choice of model architecture and training strategy will depend heavily on the specific use-case requirements, such as the need for speed, interpretability, dataset size, or deployment environment. This rich ecosystem of AI techniques underscores an exciting frontier in pest management, where technology is increasingly capable of providing timely, accurate, and scalable solutions. Figure 3 demonstrates the usage frequency of AI techniques in pest classification.
Fig 3: Usage Frequency of AI Techniques in Pest Classification
7 Benchmark Datasets
A variety of image datasets have been used in the literature for training and evaluating pest classification models. Here we describe the most prominent datasets, including their content (number of classes, images), availability, and other attributes like resolution or source, as these factors greatly influence model development and performance.
7.1 IP-102
IP102 – Insect Pest 102 dataset (Wu et al.[8]) This is one of the largest public datasets for pest recognition. It contains 102 insect pest classes across about 75,000 images. The images are collected from a variety of crops (around 8 major crop types) and include both field photographs and some lab images. The average images per class is 737, but the distribution is highly imbalance. Image resolutions vary, but many are medium resolution (e.g., 500×500 pixels or higher) in color. IP102 is publicly available (hosted with the original paper) and has become a benchmark. For example, baseline CNN accuracy was 49, highlighting its difficulty. This dataset covers a broad range of pest types (chewers, sap-suckers, borers, etc.) and is valuable for evaluating general-purpose models.
7.2 Xei-1
Xie1 (2015) – This refers to a dataset by Xie et al.[44] noted in literatur. It includes 24 common field crop insect species with a total of *1,440 images. On average 60 images per class are provided, which is quite limited for CNN training. The images likely are laboratory or catalog-style (the reference mentions insect science websites), possibly with insects photographed on plain backgrounds. Xie1 is often cited as a source of data for early pest classification research and is considered public or semi-public (used in research papers, though one might need to contact authors or sources to obtain it). The resolution is not explicitly documented, but given the nature, images may be decent resolution close-ups of insects. Due to its small size, Xie1 alone is not sufficient for deep learning, but it has been used in combination with other data or for transfer learning purposes.
7.3 Xei-2
Xie2 / D0 (2018) – Xie et al.[45] and collaborators (possibly the NBAI&R – National Bureau of Agricultural Insect Resources in China) expanded on earlier datasets. One reference indicates a dataset of 40 or 45 insect types with 4,500 images (roughly 113 images per class). This is sometimes referred to as “Xie et al. 2018” or a D0 dataset in certain paper. It includes more classes than Xie1, potentially adding stored grain pests 39 and others. This dataset is partially public; some researchers have gained access for experiment. Image resolution and format are again variable; likely a mix of lab-taken images. Xie2 serves as an intermediate scale dataset and has been used to pre-train models or evaluate alongside IP102. Its imbalance is less extreme than IP102 but still present.
7.4 Plant Village
PlantVillage is a well-known open dataset for plant disease images, containing 50,000+ images of healthy or diseased leaves for various crops. While not originally focused on insects, some categories include pest damage (e.g., pest-infested leaves). Li et al.[46] used a 39-class subset of PlantVillage that included both diseases and pest injury classe. This subset had tens of thousands of images and is public. Images are highquality, mostly lab-controlled (solid background with a single leaf). Resolution is high (many 256×256 or 512×512 after preprocessing). PlantVillage data tends to yield very high model accuracies (due to uniform images), but models trained solely on it might not generalize to field condition. Nonetheless, it’s a useful training resource for certain pests (e.g., aphid damage on leaves, etc.) and is often used for transfer learning.
7.5 Other Datasets
In addition to established benchmark datasets like IP102, PlantVillage, and LLPD, many studies rely on specialized or custom-collected datasets designed for specific pests, crops, or regional field conditions. These datasets, while often limited in size or public accessibility, provide high contextual relevance and are instrumental in developing realworld-ready AI models. For instance, Al-Shannaq et al.[3] developed a tomato pest dataset containing 859 field-captured images, augmented to 5,894 samples, which was used to train CNN classifiers on leaf damage symptoms. Berka et al.[7] introduced the CactiViT dataset—1,200 annotated images of cochineal infestations on cacti with varying resolutions—released as an open-source resource. Du et al.[13] compiled a detector-specific dataset for corn and rice pests such as borers and planthoppers, although the dataset remains unpublished. Additional smaller-scale public datasets have also been employed. Amin et al.[10] used the Kaggle Paddy Pest Dataset, which includes 648 images (135 pest-positive and 513 pest-free), for binary pest detection using hybrid quantum CNNs. The NBAI Rice Pest Dataset (Liu et al.[47], containing 5,136 images of 12 rice pest classes, was integral to early CNN-based pest classification and may have contributed to the IP102 compilation. Another recent addition is the Forestry Pest Dataset mentioned by Liu et al.[12], comprising several thousand high-resolution images across 30 insect species captured in forest environments. Although institutionally held and private, it enabled transformer-based models to achieve over 95% mAP. These datasets collectively highlight the importance of localized, context-specific data in advancing the effectiveness and deployment of AI systems in agricultural pest management. Below mentioned table 8 cover a range of crops, pest species, and image acquisition condition. Public availability and contextual fit are key for model generalizability.
Table 8: Summary of Notable Pest Image Datasets.
Ref Dataset (Year) Classes Images Public Typical Image Resolution Notes
[8] IP102 (2019) 102 $\sim$75,000 Yes $\times500$ px (varied; many field shots) Benchmark for fine-grained pest recognition.
[44] Xie1 (2015) 24 1,440 Semi $\sim$$\times600$ px (est., lab images) Early dataset of common pest; small per class (60). Often used for initial model tests.
[45] Xie2/D0 (2018) $\sim$40–45 $\sim$4,500 Semi $\sim$$\times800$ px (est.) Includes more species (possibly storage pests).
[46] PlantVillage-39 (2016) 39 (mixed) $\sim$54,000 (subset) Yes $\times256$ px (processed) Lab-like images, high quality.
[3] Tomato Pest (2025) 2 (infested vs healthy) 859 (raw) / 5,894 (aug) No $\times1080$ px (smartphone) Augmented for training. Focused binary classification.
[…]
[47] Rice Pest 12 (2016) 12 5,136 No $\sim$$\times400$ to $\times800$ px (field) Rice field pest image. Moderately sized, used in early CNN work.
[13] Corn-Rice Pest (2024) $\sim$20 (est.) few thousand (est.) No $\times1024$ px (drone/field photos) Du et al.’s YOLOv7-PSAFP data. Focused on small pests in field. Possibly will be released in future.
[…]
8 Challenges
8.1 Dataset Imbalance
A foundational challenge in pest classification lies in the issue of imbalanced and in- sufficient datasets. Most existing datasets show a stark imbalance in class distribution common pests tend to dominate with thousands of images, while rare or region- specific pests may have only a handful of annotated samples. This imbalance can significantly skew model performance, often resulting in high accuracy for dominant classes and poor generalization for underrepresented ones. Fu et al.[35] directly noted this issue, reporting reduced model 41 accuracy for classes with fewer samples. To counteract this, researchers have adopted a range of strategies. Data augmentation is ubiquitous, used to synthetically expand dataset diversity. Loss function modifications like focal loss (Venkatasaichandrakanth Iyapparaja,[15]) aim to reduce the bias towards frequent classes. However, these methods offer only partial solutions. The deeper issue is the limited availability of comprehensive, high-quality pest image datasets. Al-Shannaq et al.[3] and Lin et al.[20] both identified the scarcity of annotated images as a critical bottleneck for robust AI development. Innovative responses are emerging. Kar et al.[24] demonstrated how selfsupervised learning particularly the BYOL approach could leverage large volumes of unlabeled data to pre-train models that approach supervised performance. Moving forward, promising directions include the creation of openaccess, crowdsourced pest image repositories and the use of few-shot and semisupervised learning techniques. These approaches could help democratize data availability, enabling more inclusive and generalized AI systems for pest classification. Table 9 highlights the challenges and strategies for handling class imbalance and data scarcity in agricultural pest recognition.
Table 9: Key Studies Addressing Data Scarcity and Long-Tail Distribution in Pest Recognition.
Reference Year Description
[35] 2024 Noted significant accuracy drop on under-represented pest classes; highlighted how long-tailed training data skewed ViT performance.
[15] 2024 Applied focal loss to reduce bias toward frequent classes in pest classification, giving more weight to minority class examples.
[3] 2025 Emphasized the critical shortage of localized pest image datasets (e.g., for Jordan crops), identifying data scarcity as a key limitation.
[20] 2023 Similarly pointed out the lack of comprehensive, annotated pest image repositories; called for community efforts in data collection and sharing.
[24] 2022 Demonstrated self-supervised pre-training (BYOL) on unlabeled pest images to combat data scarcity; achieved near supervised-level accuracy with far fewer labeled examples.
[8] 2019 Released IP102, a 75,000+ image dataset of 102 pest species with a naturally imbalanced distribution, a benchmark that spurred research into long-tail learning for pests.
[49] 2021 Created AgriPest, a large-scale pest dataset ($\sim$49.7k images, 14 species) captured in real field conditions; aims to provide a realistic benchmark for training and evaluation.
8.2 Complex Field Environments
Another major hurdle is the variability and complexity of real-world field conditions. Unlike controlled laboratory environments, crop fields present cluttered backgrounds composed of leaves, stems, soil, and debris, combined with inconsistent lighting conditions (e.g., glare, shadows, low-light scenarios). These elements introduce noise and visual distractions that significantly challenge model robustness and generalization. This challenge has been widely acknowledged in the literature. Liu et al.[12] introduced a novel Region-based Pyramid Self-Attention (RPSA) mechanism to specifically account for background context, demonstrating improved detection performance in visually busy field images. Du et al.[13] tackled the issue by emphasizing hard ex- ample mining during training, which helped reduce false positives arising from back- ground confusion. Despite these advancements, the problem is far from solved. Wang et al.[6] pointed out that their model’s accuracy still degraded under challenging environmental conditions like rain, lens fog, or strong shadows. Such findings point to a need for improved data augmentation strategies that simulate various lighting and environmental conditions. Additionally, domain adaptation techniques where models are trained across diverse domains or geographical settings—could improve generalization and robustness. Future work must continue to address the environmental variability inherent in realworld deployment scenarios.
Table 10: Selected Studies Tackling Field Complexity and Robustness in Pest Detection.
Reference Year Description
[12] 2025 Introduced a Region-based Pyramid Self-Attention in their detector to handle cluttered backgrounds; improved pest detection in visually busy field images by focusing on relevant regions.
[13] 2024 Employed hard example mining during training to teach the model using cluttered, difficult images; reduced false positives from leaves and background noise.
[17] 2025 Observed model accuracy degrading with rain, fog, or harsh shadows; highlighted need for augmentation simulating weather and lighting variability for robust performance.
[41] 2022 Developed a deformable residual network with a global context module for pest detection in complex scenes; achieved 77.8% mAP on 21 pest classes in field conditions by explicitly modeling context.
[39] 2023 Proposed environment-invariant training techniques (e.g., multi-domain data and adaptive normalization) enabling their Insect-YOLO to maintain accuracy across different backgrounds and lighting conditions.
8.3 Detection of Small Pests
A significant technical difficulty lies in the detection of small pests, which often occupy only a tiny portion of the image. This creates problems for both classification and detection models, as the visual features of small objects are often lost in deep layers of neural networks. It’s a known limitation even in general computer vision, and it’s especially acute in agricultural applications where pests can be millimeters in size. Several researchers have proposed model-specific solutions. Hu et al.[25] added an additional fine-scale detection head to YOLO to capture small pest signatures more effectively. Wang et al.[6] incorporated Convolutional Block Attention Module (CBAM) attention into their detection pipeline, enhancing sensitivity to small, localized objects. Likewise, the systematic review by Teixeira et al.[9] identified small object detection as a consistent challenge across detection benchmarks. Potential future directions include leveraging super-resolution techniques to enhance small objects before detection or combining traditional RGB imaging with specialized sensors (macro lenses, hyperspectral cameras). Additionally, techniques like multi-scale feature pyramids and transformer-based aggregation (such as in Liu et al.[12] have proven valuable in maintaining detailed spatial information. At the hard- ware level, deploying high-resolution, close-range cameras such as smart traps or crop-mounted sensors—can support better detection, though they require corresponding algorithmic adjustments to manage the increased data load.The below mentioned studies in the table 11 demonstrate state of the art techniques for improving the detection and classification of tiny insects in real and lab.
Table 11: Selected Studies Targeting Small Pest Detection in Agricultural Images.
Reference Year Description
[25] 2023 Added an extra fine-scale detection head in their YOLO-based model to capture very small pests; improved detection of tiny insects that standard heads missed.
[6] 2025 Incorporated CBAM attention into the detection pipeline, heightening the model’s sensitivity to small, localized pest features (e.g., aphids on a leaf).
[9] 2023 Identified small object detection as a major challenge across studies; noted that models perform significantly worse on insects that occupy few pixels, urging focus on multi-scale methods.
[50] 2024 Developed a method for detecting whiteflies on yellow sticky traps (small targets); used high-resolution input and an improved YOLOv5, achieving high precision on pests $\sim$3mm in size.
[51] 2023 Applied image super-resolution on suspected pest regions before classification; demonstrated that upsampling tiny pest images improved identification accuracy in a lab setting.
8.4 Computational Cost
Beyond accuracy, the real-world deployment of pest detection models demands computational efficiency and responsiveness. Several studies have highlighted the tension between model complexity and deployability. High-capacity models like Vision Trans- formers offer superior accuracy but require considerable resources, making them impractical for edge devices or mobile applications commonly used in agricultural fields. Al-Shannaq et al. (2025) articulated this challenge explicitly, targeting mobile deployment as a core use case. Similarly, Liu et al.[12] and Saranya et al.[37] focused on developing models that could operate under edge constraints. Berka et al.[7] successfully deployed a pest classification system on a mobile device, showing that real-time applications are feasible but require careful model design. To reduce computational demand, several strategies are being explored. Xiao et al.[23] used GA-GhostNet, a streamlined CNN architecture that balances efficiency and accuracy by significantly reducing model parameters. Other potential approaches include pruning, quantization, and knowledge distillation where a small “student” model is trained to mimic the behavior of a large, accurate “teacher” model. Although knowledge distillation remains underexplored in pest classification, it holds promise as a technique to deliver high performance in resource-constrained environments. the table 12 highlights the studies progress and innovation in real time, resource efficient pest identification for mobile and edge computing.
Table 12: Recent Studies on Mobile/Edge Deployment for Pest Detection.
Reference Year Description
[3] 2025 Emphasized mobile deployment requirements; aimed to develop models that run on smartphones for on-site pest identification.
[7] 2023 Demonstrated a successful on-device pest classification app; used model optimization (quantization, etc.) to achieve real-time inference on mobile hardware.
[23] 2024 Utilized GA-GhostNet (an efficient CNN) to balance accuracy and speed; significantly reduced model parameters while maintaining high precision.
[37] 2024 Designed their hybrid model with edge constraints in mind (e.g., limiting extra parameters from HP-MHA module) to ensure feasibility of deployment.
[52] 2022 Created a lightweight YOLOv3-based detector ($\sim$5M params) that runs $\sim$20 FPS on edge device; achieved >>80% mAP, illustrating a strong accuracy-speed trade-off for field use.
[53] 2024 Improved YOLOv8 architecture for efficiency (small dataset of 7 classes); obtained 82.3% mAP with only 2M parameters, suitable for low-power deployment.
[54] 2015 Proposed knowledge distillation, a method yet to be fully explored in pest classification, where a smaller model learns from a larger model’s outputs, potentially yielding a compact, accurate pest classifier.
8.5 Limited Generalization
Another open issue is the limited generalization capacity of current models. Many are trained and evaluated on narrowly defined datasets often from a single region, crop type, or image source. While they perform well within these boundaries, their accuracy can degrade sharply when applied to new environments, crops, or pest varieties. This poses a major barrier to the widespread adoption of AI in diverse agricultural settings. Some progress has been made. Xiao et al.[23] explored cross-dataset transfer, showing that a model trained on pest images could be fine-tuned for disease recognition. However, transfer learning across different pest datasets or geographic regions is still largely under-investigated. GA-GhostNet’s limitations in generalization further underscore this gap. Lin et al.[20] even speculated on broader applications, including predictive modeling of pest outbreaks based on environmental data, though that remains speculative at this stage. A future-ready pest detection system must be able to handle new classes and unforeseen pest appearances. This calls for approaches like open-set recognition, continual learning, and active learning. In particular, models that can flag uncertain pre- dictions for human labeling and then incorporate new knowledge over time would significantly improve adaptability. This is increasingly important as climate change alters pest distributions, introducing novel species into previously unaffected areas. Therefore, model adaptability and long-term learning capabilities will be essential for sustainable pest management in evolving agricultural ecosystems. The studies mentioned in the table 13 illustrate the current best practices and emerging strategies for robust, adaptable pest and plant disease AI across varied and evolving agricultural contexts.
Table 13: Recent Advances and Challenges in Generalization, Transfer Learning, and Adaptability for Pest and Plant Disease AI.
Reference Year Description
[23] 2024 Found that models trained on one pest dataset did not generalize well to another; cross-dataset transfer learning required fine-tuning, indicating domain-specific bias.
[20] 2023 Highlighted the need for broader generalization; suggested integrating environmental data for predictive modeling, but noted this remains speculative without further research.
[42] 2022 Noted that transfer learning and domain adaptation are crucial for adapting models trained in one context to new settings with minimal retraining.
[19] 2023 Demonstrated that out-of-distribution detection can allow pest classifiers to abstain on unknown species, improving reliability and enabling open-set recognition.
[55] 2019 Reviewed continual learning techniques; relevant to pest AI for updating models with new pest classes over time without forgetting old ones, thereby enhancing long-term generalization.
[56] 2022 Applied cross-domain few-shot learning for plant disease identification; concept can be extended to pests, enabling models to learn new pest categories from only a few examples, enhancing adaptability.
9 Conclusion
This review discusses how pest classification has quickly developed with deep learning, from conventional CNNs to stronger hybrid and transformer-based models that yield superior accuracy and versatility in agricultural settings. After reviewing 37 studies, it is apparent that the majority of work involves crop plants such as rice, groundnut, and tomato, with emerging interest in others like cactus and tea. Whereas image based approaches utilizing platforms such as CCTV and drones are prominent, sensor based and smartphone friendly solutions are making inroads for realtime application in the field. Nonetheless, significant bottlenecks exist, such as skewed datasets, difficulty seeing small or concealed pests, and models that don’t generalize well between regions or crops. There are promising new improvements in quantum inspired CNNs, data augmentation techniques such as diffusion models, and edge friendly versions of YOLO. Future work must address building larger and diverse open datasets, developing energy efficient models that can be deployed in the field, and building systems that generalize well under varying agricultural conditions. Transitioning research to farms will necessitate strong collaboration among AI researchers and agronomists.